OpenStudy (anonymous):

Translate the sentence into an equation. Five times the difference of a number and 8 is 45. A. 5 ∙ (m – 8) = 45 B. 5m – 8 = 45 C. m – 8 ∙ 5 = 45 D. 5 ∙ (m + 8) = 45

OpenStudy (jojokiw3):

Try A.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Zach's baseball team won 32 games last season. They played a total of 45 games. How many games did they lose overall? Let l represent the number of games that Zach's team lost. A. l – 32 = 45 B. l + 32 = 45 C. l + 45 = 32 D. 32l = 45